# The stability of propagating plane inertial waves in rotating fluids

### v1

@@ Abstract
Inertial waves transport energy and momentum in rotating fluids and are a major contributor to mixing and tidal dissipation in Earth’s oceans, gaseous planets, and stellar interiors. However, their stability and breakdown mechanisms are not fully understood. We examine the linear stability and nonlinear breakdown of finite-amplitude propagating plane inertial waves using Floquet theory and direct numerical simulations. The Floquet analysis generalizes previous studies as it is valid for arbitrary perturbation wavelengths and primary wave amplitudes. We find that the wavenumber orientation of the most unstable perturbations depends strongly on the wave frequency and weakly on the wave amplitude. The most unstable perturbations have wavelengths that are small relative to the primary wave wavelength for low wave amplitudes, but become comparable for large wave amplitudes. We then use direct numerical simulations to follow the nonlinear breakdown of the wave and examine how the wave energy is either dissipated in a forward cascade or accumulated into long-lived geostrophic modes. Simulations reveal that the conversion efficiency into geostrophic modes increases with increasing wave amplitude, as expected for pumping of geostrophic modes by nearly-resonant triadic interactions. We also find that the conversion efficiency increases with decreasing primary wave frequency, which may be due to the more efficient coupling of quasi-2D waves to geostrophic modes. These results on the stability and breakdown of single plane inertial waves provides additional foundation for understanding the role of inertial waves in rotating turbulence, transport properties of inertial wave beams, and inertial wave propagation in more complex environments such as those with magnetic fields or shear flows.
@@ I Introduction
Rotating fluids support internal waves known as inertial waves (IWs), which result from the interaction between fluid inertia and the Coriolis force [1]. In a fluid with a rotation rate $\Omega>0$, IWs are transverse, dispersive waves with frequencies $|\omega|\leq 2\Omega$ determined only by the angle of propagation relative to the rotation axis. They are fundamental in rotating flows because they mediate turbulent mixing and non-locally transport angular momentum and energy [2, 3, 4, 5, 6], thereby affecting long-term dynamical evolution of rotating fluid bodies.
@@
The excitation, propagation, and dissipation of IWs has been studied in many natural systems, including stars, planets, and moons. Tidally excited IWs in star-planet systems can meditate angular momentum exchange as they propagate in convective cavities and dissipate through wave-wave or wave-shear flow instabilities [7, 8, 9]. In planet-moon systems, dissipation of tidally-driven IWs can provide significant internal heating [10], in addition to affecting their orbital and rotational evolution [e.g. 11, 12, 13, 14, 15, 16, 17]. In terrestrial contexts, near-inertial waves in the upper oceans are a major contributor to the turbulent mixing that controls the depth of the oceanic mixed layer, impacting atmosphere-ocean coupling and the climate [18]. Observational studies have successfully identified IWs in several natural systems, including the solar convective envelope [19, 20, 21, 22], Earth’s outer core [e.g. 23], and Earth’s oceans [e.g. 24, 25, 26].
@@
Although IWs are important in geophysical and astrophysical applications, they have been less well studied than internal gravity waves (IGWs), which propagate in stably stratified fluids and also contribute to mixing in Earth’s oceans [e.g. 2] and stellar interiors [e.g. 27]. Fortunately, many of the techniques used to study IGWs can be applied to IWs because these waves represent different limits of the same internal wave branch in a rotating and stratified incompressible fluid. As a result, they share many properties and, in particular, both are subject to triadic resonance instabilities (TRI) of small-amplitude waves, which transfer energy and momentum from a “primary wave” to pairs of “daughter waves” that satisfy temporal and spatial resonance conditions. Early studies of IGW stability focused on a subset of TRI known as parametric subharmonic instabilities (PSI) where the daughter wave frequencies are roughly half of the primary wave frequency [for plane propagating IGWs, e.g. 28, 29, 30, 31]. One of the most comprehensive studies of IGW stability employed a numerical Floquet method applicable for any finite amplitude of the primary wave [32]. This approach also allows exploration of the stability of perturbations with wavelengths comparable to (or larger than) the primary wave, rather than being restricted to infinitesimally small wavelengths.
@@
Some similar efforts have been made to analyze the stability of IWs. Early work showed that standing IWs are generically unstable in both Cartesian [33, 34] and cylindrical geometries [35]. IW attractors in 2D geometries have also been studied experimentally, numerically, and theoretically [36, 37], where unstable perturbations were restricted to the plane of the rotation axis and the attractor. Relaxing this 2D assumption, Brunet et al. [38] found that IW attractors are broadly unstable to 3D modes. Recently, the first analytical and experimental stability analysis of plane propagating IWs by Mora et al. [39] found that the TRI for small-amplitude propagating IWs is also generally 3D—unstable modes efficiently redistribute energy into the horizontal plane normal to the rotation axis. Most recently, Abhiram and Mathur [40] extended the linear stability analysis to finite-amplitude waves using a local stability approach, characterizing the amplitude dependence of infinitesimally short-wavelength perturbations.
@@
In this work, we extend the linear stability analysis of finite-amplitude, plane, propagating IWs to arbitrary wavelength perturbations and examine the nonlinear breakdown of IW. We utilize the Floquet method of Lombard and Riley [32] for the stability analysis in Section II and supplement with asymptotic analysis using the method of multiple timescales [e.g. 11, 41, 42, 43, Appendix B]. We then use direct numerical simulations (DNS) of propagating IWs in Section III to cross-validate the Floquet theory and examine the ultimate nonlinear breakdown of the wave, with a focus on the conversion efficiency of the wave energy into geostrophic modes.
@@ II.1 Model and setup for the linear stability analysis
$\bm{U}=\mathrm{Re}[A\bm{h}_{s}e^{i(\bm{k}\cdot\bm{x}-\omega t))}],\quad\bm{h}_{s}=\hat{\bm{k}}\times\hat{\bm{\Psi}}+is\hat{\bm{\Psi}},\quad\hat{\bm{\Psi}}=\frac{\hat{\bm{k}}\times\hat{\bm{z}}}{|\hat{\bm{k}}\times\hat{\bm{z}}|},\quad s=\pm 1,$ (2)
where $A$ is the constant amplitude of the wave and $\hat{\bm{k}}$ is the unit vector in the direction of the wavevector $\bm{k}$ (with magnitude $k=|\bm{k}|$). One can see that individual wave solutions have a vanishing nonlinear advection term, $\bm{U}\cdot\nabla\bm{U}=0$, and hence are exact solutions for any amplitude. The real frequency $\omega$ satisfies the dispersion relation
@@
We are interested in the stability of plane IWs, viewing their flow (and pressure) field as a spatially and temporally periodic base flow. Linear stability analysis is most convenient in a frame translating with the phase velocity of the wave. By rotating the vertical axis to align with $\hat{\bm{k}}$ and applying a velocity boost along $\hat{\bm{k}}$ with the phase speed $\omega/k$, the base flow transforms from a periodic function of space and time coordinates in the rotating frame, $\bm{U}=\bm{U}(\bm{k}\cdot\bm{x}-\omega t)$, to a periodic function of a single “phase” variable $\phi=\bm{k}\cdot\bm{x}-\omega t$ in the comoving frame, $\bm{U}^{\prime}=\bm{U}^{\prime}(\phi)$.
The explicit transformation to the comoving frame is carried out as follows. Without loss of generality, consider a primary wave with a wavevector $\bm{k}=(k_{x},0,k_{z})$ in the $x$-$z$ plane, and with a real amplitude $A$. The units for space and time in the comoving coordinates are chosen based on the wavenumber and frequency of the primary wave. The transformation from the rotating frame coordinates $(x,y,z)$ to the comoving dimensionless coordinates $(x^{\prime},y^{\prime},\phi)$ (with unit vectors $\hat{\bm{x^{\prime}}}$, $\hat{\bm{y}^{\prime}}$, and $\hat{\bm{\phi}}$) is then
@@
$u^{\prime}=\frac{1}{|\omega|}\left(k_{z}u-k_{x}w\right),\;v^{\prime}=\frac{k}{|\omega|}v,\;w^{\prime}=\frac{1}{|\omega|}\left(k_{x}u+k_{z}w\right)-\frac{\omega}{|\omega|},\;p^{\prime}=\frac{k^{2}}{\omega^{2}}p.$ (5)
@@
The new dimensionless amplitude $A^{\prime}$ (which we take to be real and positive) is essentially a Rossby number of the wave, quantifying the ratio of the advective rate $\sim A\,k$ to the wave frequency $\omega$. The horizontal components of the velocity field ($U^{\prime},V^{\prime}$) trace out a right/left-handed corkscrew along $\hat{\phi}$ for $s=\pm 1$, while the vertical component corresponds to a uniform motion along the $\hat{\phi}$ axis. Viscous effects on the primary wave have again been ignored (the limit $\rm Ek\rightarrow 0$, and hence $\omega\in\mathbb{R}$) because we are interested in the regime in which the growth rates of perturbations considerably exceed the viscous damping rate of the primary wave. Hereafter, viscous effects on the primary wave are neglected, which also enables the use of Floquet theory for the linear stability analysis.
We now consider the evolution of perturbations $\bm{\widetilde{u}}^{\prime}$ with small amplitudes $|\bm{\widetilde{u}}^{\prime}|\ll|A^{\prime}|$ on top of the base flow $\bm{U}^{\prime}$. Substituting $\bm{u}^{\prime}=\bm{U}^{\prime}+\bm{\widetilde{u}}^{\prime}$ and dropping quadratic terms $O(|\bm{\widetilde{u}}^{\prime}|^{2})$ gives the governing equations at first order for the perturbations,
@@
$\widetilde{f}^{\prime}=\mathrm{Re}\left[f^{\dagger}(\phi)e^{i(\alpha x^{\prime}+\beta y^{\prime})+\sigma t^{\prime}}\right],$ (9)
@@
where $\widetilde{f}^{\prime}$ is one of $\widetilde{u}^{\prime}$, $\widetilde{v}^{\prime}$, $\widetilde{w}^{\prime}$, $\widetilde{p}\,^{\prime}$, and $f^{\dagger}$ is to be determined. The horizontal wavenumbers $\alpha$ and $\beta$ are real while the frequency $\sigma$ is complex in general. Unstable modes are those with $\mathrm{Re}[\sigma]>0$. Substitution of Equation 9 reduces the system to a set of coupled ordinary differential equations in $\phi$:
$(S-\mathrm{Ek}L_{2})u^{\dagger}+w^{\dagger}\frac{\mathrm{d}U^{\prime}}{\mathrm{d}\phi}-\frac{k_{z}}{|k_{z}|}v^{\dagger}$ $=-i\alpha p^{\dagger},$ (10)
$(S-\mathrm{Ek}L_{2})v^{\dagger}+w^{\dagger}\frac{\mathrm{d}V^{\prime}}{\mathrm{d}\phi}+\frac{k_{z}}{|k_{z}|}u^{\dagger}+\frac{k_{x}}{|k_{z}|}w^{\dagger}$ $=-i\beta p^{\dagger},$ (11)
$(S-\mathrm{Ek}L_{2})w^{\dagger}-\frac{k_{x}}{|k_{z}|}v^{\dagger}$ $=-\frac{\mathrm{d}p^{\dagger}}{\mathrm{d}\phi},$ (12)
$i\alpha u^{\dagger}+i\beta v^{\dagger}+\frac{\mathrm{d}w^{\dagger}}{\mathrm{d}\phi}$ $=0,$ (13)
@@
where $S=\sigma+i\alpha U^{\prime}+i\beta V^{\prime}+W^{\prime}\frac{\mathrm{d}}{\mathrm{d}\phi}$, $L_{2}=-\alpha^{2}-\beta^{2}+(\mathrm{i}\gamma+\frac{\mathrm{d}}{\mathrm{d}\phi})^{2}$, and the last equation is the incompressibility constraint. The coefficients $U^{\prime}$ and $V^{\prime}$ are $2\pi$-periodic, so Floquet theory implies that $f^{\dagger}$ takes the form
$f^{\dagger}(\phi)=e^{i\gamma\phi}\hat{f}(\phi)=e^{i\gamma\phi}\sum_{n=-\infty}^{n=\infty}f^{\prime}_{n}e^{in\phi}.$ (14)
@@
where $\hat{f}(\phi)$ is a $2\pi$-periodic function that has been expanded in Fourier modes with constant complex coefficients $f^{\prime}_{n}$. Floquet theory allows $f^{\dagger}$ to be aperiodic since $\gamma$ need not take an integer value. However, $\gamma$ must be real for physical solutions that are bounded along $\phi$. Substitution of the Floquet ansatz (Equations 9 and 14) and the base flow (Equation 7) into the governing equations for the perturbations (Equation 8) reduces the problem to a set of coupled algebraic equations. Dropping the prime notation for perturbations going forward, we obtain
@@
$(\sigma-\mathrm{Ek}L_{2n})u_{n}+\frac{i\alpha A^{\prime}}{2}(u_{n+1}+u_{n-1})-\frac{\beta sA^{\prime}}{2}(u_{n+1}-u_{n-1})$- $\frac{i\omega}{|\omega|}(\gamma+n)u_{n}-\frac{iA^{\prime}}{2}(w_{n+1}-w_{n-1})$
$-\frac{k_{z}}{|k_{z}|}v_{n}=-i\alpha p_{n},$ (15)
@@
$(\sigma-\mathrm{Ek}L_{2n})v_{n}+\frac{i\alpha A^{\prime}}{2}(v_{n+1}+v_{n-1})-\frac{\beta sA^{\prime}}{2}(v_{n+1}-v_{n-1})$- $\frac{i\omega}{|\omega|}(\gamma+n)v_{n}+\frac{sA^{\prime}}{2}(w_{n+1}+w_{n-1})$
$+\frac{k_{z}}{|k_{z}|}u_{n}+\frac{k_{x}}{|k_{z}|}w_{n}=-i\beta p_{n},$ (16)
@@
$(\sigma-\mathrm{Ek}L_{2n})w_{n}+\frac{i\alpha A^{\prime}}{2}(w_{n+1}+w_{n-1})-\frac{\beta sA^{\prime}}{2}(w_{n+1}-w_{n-1})$- $\frac{i\omega}{|\omega|}(\gamma+n)w_{n}-\frac{k_{x}}{|k_{z}|}v_{n}=-i(\gamma+n)p_{n},$ (17)
$i\alpha u_{n}+i\beta v_{n}+i(\gamma+n)w_{n}$ $=0,$ (18)
@@
where $L_{2n}\equiv-\alpha^{2}-\beta^{2}-(\gamma+n)^{2}$. The linear stability of the wave can now be determined by solving this algebraic eigenvalue problem for the complex frequency $\sigma$ as a function of the parameters of the primary wave $\{A^{\prime},\omega,s\}$, the horizontal wavenumbers of the perturbations in the comoving frame $\{\alpha,\;\beta\}$, and the Floquet parameter $\gamma$. For completeness, the energy conservation equation for this Floquet system is derived in Appendix A.
@@
The eigenvalue problem given by Equations II.118 has several symmetries that reduce the parameter space over which to search for unstable modes. Several symmetries are shared with the Floquet system for IGWs [32], which has three symmetries: 1) symmetry for eigenvectors $\bm{V}_{n}\equiv[u_{n},v_{n},w_{n},p_{n}]\to\bm{V}_{n}^{*}$ (where * denotes the complex conjugate and $\bm{V}^{*}_{n}=\bm{V}_{-n}$), $(\alpha,\beta,\gamma)\to(-\alpha,-\beta,-\gamma)$, and $\sigma\to\sigma^{*}$, 2) invariance of $\sigma$ to integer shifts in the Floquet parameter $\gamma\rightarrow\gamma+1$, 3) invariance of $\sigma$ in the $\beta$-plane, $\beta\rightarrow-\beta$ and $v_{n}\rightarrow-v_{n}$. The Floquet system for IWs satisfies the first two symmetries, but lacks the third because the non-zero $V^{\prime}\neq 0$ component of the base flow interacts with perturbations that have $\beta\neq 0$ (unlike for IGWs where $V^{\prime}=0$). As a result, the Floquet system for IWs must be explored in the upper (or lower) half of the $\alpha-\beta$ plane and for all $0\leq\gamma<1$.

### v2

@@ Abstract
Inertial waves transport energy and momentum in rotating fluids and are a major contributor to mixing and tidal dissipation in Earth’s oceans, gaseous planets, and stellar interiors. However, their stability and breakdown mechanisms are not fully understood. We examine the linear stability and nonlinear breakdown of finite-amplitude propagating plane inertial waves using Floquet theory and direct numerical simulations. The Floquet analysis generalizes previous studies as it is valid for arbitrary perturbation wavelengths and primary wave amplitudes. We find that the wave vector orientation of the most unstable perturbations depends strongly on the wave frequency and weakly on the wave amplitude. The most unstable perturbations have wavelengths that are small relative to the primary wave wavelength for low wave amplitudes, but become comparable for large wave amplitudes. We then use direct numerical simulations to follow the nonlinear breakdown of the wave and examine how the wave energy is either dissipated in a forward cascade or accumulated into long-lived geostrophic modes. Simulations reveal that the conversion efficiency into geostrophic modes increases with increasing wave amplitude, as expected for pumping of geostrophic modes by nearly-resonant triadic interactions. We also find that the conversion efficiency increases with decreasing primary wave frequency, which may be due to the more efficient coupling of quasi-2D waves to geostrophic modes. These results on the stability and breakdown of single plane inertial waves provide an additional foundation for understanding the role of inertial waves in rotating turbulence, transport properties of inertial wave beams, and inertial wave propagation in more complex environments such as those with magnetic fields or shear flows.
@@ I Introduction
Rotating fluids support internal waves known as inertial waves (IWs), which result from the interaction between fluid inertia and the Coriolis force [34]. In a fluid with a rotation rate $\Omega>0$, IWs are transverse, dispersive waves with frequencies $|\omega|\leq 2\Omega$ determined only by the angle of propagation relative to the rotation axis. They are fundamental in rotating flows because they mediate turbulent mixing and non-locally transport angular momentum and energy [62, 25, 32, 3, 38], thereby affecting the long-term dynamical evolution of rotating fluid bodies.
@@
The excitation, propagation, and dissipation of IWs have been studied in many natural systems, including stars, planets, and moons. Tidally excited IWs in star-planet systems can mediate angular momentum exchange as they propagate in convective zones and dissipate through wave-wave or wave-shear flow instabilities [29, 5, 6]. In planet-moon systems, dissipation of tidally-driven IWs can provide significant internal heating [58], in addition to affecting their orbital and rotational evolution [e.g. 41, 54, 55, 16, 44, 9, 13]. In terrestrial contexts, near-inertial waves in the upper oceans are a major contributor to the turbulent mixing that controls the depth of the oceanic mixed layer, impacting atmosphere-ocean coupling and the climate [37]. Observational studies have successfully identified IWs in several natural systems, including the solar convective envelope [48, 31, 14, 35], Earth’s outer core [e.g. 2], and Earth’s oceans [e.g. 30, 36, 63].
@@
Although IWs are important in geophysical and astrophysical applications, they have been less well studied than internal gravity waves (IGWs), which propagate in stably stratified fluids and also contribute to mixing in Earth’s oceans [e.g. 62] and stellar interiors [e.g. 11]. These two wave types, however, derive from the same internal wave branch in a rotating and stratified incompressible fluid. As a result, they share many properties and, in particular, both are subject to triadic resonance instabilities (TRI) of small-amplitude waves, which transfer energy and momentum from a “primary wave” to pairs of “daughter waves” that satisfy temporal and spatial resonance conditions. Early studies of IGW stability focused on a subset of TRI known as parametric subharmonic instabilities (PSI) where the daughter wave frequencies are roughly half of the primary wave frequency [for plane propagating IGWs, e.g. 50, 28, 42, 61]. One of the most comprehensive studies of IGW stability employed a numerical Floquet method applicable for any finite amplitude of the primary wave [47]. This approach also allows exploration of the stability of perturbations with wavelengths comparable to (or larger than) the primary wave, rather than being restricted to infinitesimally small wavelengths.
@@
Similar efforts have been made to analyze the stability of IWs. Early work showed that standing IWs are generically unstable in both Cartesian [46, 51] and cylindrical geometries [40]. IW attractors in 2D geometries have also been studied experimentally, numerically, and theoretically [17, 39], where unstable perturbations were restricted to the plane of the rotation axis and the attractor. Relaxing this 2D assumption, Brunet et al. [20] found that IW attractors are broadly unstable to 3D modes. For plane propagating IWs, the first analytical and experimental stability analysis by Mora et al. [52] found that the TRI for small-amplitude primary waves is also generally 3D—unstable modes efficiently redistribute energy into the horizontal plane normal to the rotation axis. Recently, Abhiram and Mathur [1] extended the linear stability analysis to finite-amplitude waves using a local stability approach appropriate for infinitesimally short-wavelength perturbations, characterizing the dependence of the most unstable perturbations on the primary wave amplitude and frequency.
@@
In this work, we extend the linear stability analysis of finite-amplitude, plane, propagating IWs to arbitrary wavelength perturbations and examine the nonlinear breakdown of IWs. We utilize the Floquet method of Lombard and Riley [47] for the stability analysis in Section II and supplement this with asymptotic analysis using the method of multiple timescales [e.g. 41, 56, 8, 23, Appendix B]. We then use direct numerical simulations (DNS) of propagating IWs in Section III to cross-validate the Floquet theory and examine the ultimate nonlinear breakdown of the wave, with a focus on the conversion efficiency of the wave energy into geostrophic modes.
@@ II.1 Model and setup for the linear stability analysis
$\bm{U}=\mathrm{Re}[A\bm{h}_{s}e^{\mathrm{i}(\bm{k}\cdot\bm{x}-\omega t))}],\quad\bm{h}_{s}=\hat{\bm{k}}\times\hat{\bm{\Psi}}+\mathrm{i}s\hat{\bm{\Psi}},\quad\hat{\bm{\Psi}}=\frac{\hat{\bm{k}}\times\hat{\bm{z}}}{|\hat{\bm{k}}\times\hat{\bm{z}}|},\quad s=\pm 1,$ (2)
where $A$ is the amplitude of the wave and $\hat{\bm{k}}$ is the unit vector in the direction of the wavevector $\bm{k}$ (with magnitude $k=|\bm{k}|$). One can see that individual wave solutions have a vanishing nonlinear advection term, $\bm{U}\cdot\nabla\bm{U}=0$, and hence are exact solutions for any amplitude. The real frequency $\omega$ satisfies the dispersion relation
@@
The linear stability analysis of a plane IW is most convenient in a frame translating with the wave phase velocity. By rotating the coordinate frame to align $\hat{\bm{z}}$ with $\hat{\bm{k}}$ and applying a velocity boost along $\hat{\bm{z}}$ with the phase speed $\omega/k$, the base flow transforms from a periodic function of space and time coordinates in the rotating frame, $\bm{U}=\bm{U}(\bm{k}\cdot\bm{x}-\omega t)$, to a periodic function of a single spatial variable $\phi$ (the wave phase) in the comoving frame, $\bm{U}^{\prime}=\bm{U}^{\prime}(\phi)$.
The explicit transformation to the comoving frame is carried out as follows. Without loss of generality, consider a primary wave with a wavevector $\bm{k}=(k_{x},0,k_{z})$ in the $x$-$z$ plane, with $k_{x}\geq 0$ and a real amplitude $A>0$. The units for space and time in the comoving coordinates are chosen based on the wavenumber and frequency of the primary wave. The transformation from the rotating frame coordinates $(x,y,z)$ to the comoving dimensionless coordinates $(x^{\prime},y^{\prime},\phi)$ (with unit vectors $\hat{\bm{x^{\prime}}}$, $\hat{\bm{y}^{\prime}}$, and $\hat{\bm{\phi}}$) is then
@@
$u^{\prime}=\frac{1}{|\omega|}\left(k_{z}u-k_{x}w\right),\;v^{\prime}=\frac{k}{|\omega|}v,\;w^{\prime}=\frac{1}{|\omega|}\left(k_{x}u+k_{z}w\right)-\frac{\omega}{|\omega|},\;p^{\prime}=\frac{k^{2}}{\rho\omega^{2}}p.$ (5)
@@
The new dimensionless amplitude $A^{\prime}$ is essentially a Rossby number of the wave, quantifying the ratio of the advective rate $\sim A\,k$ to the wave frequency $|\omega|$. The horizontal components of the velocity ($U^{\prime},V^{\prime}$) trace out a right/left-handed corkscrew along $\hat{\phi}$ for $s=\pm 1$, while the vertical component corresponds to uniform motion along the $\hat{\phi}$ axis.
We now consider the stability of perturbations $\bm{\widetilde{u}}^{\prime}$ with small amplitudes $|\bm{\widetilde{u}}^{\prime}|\ll|A^{\prime}|$ on top of the inviscid base flow $\bm{U}^{\prime}$. We neglect viscous damping of the primary wave because we are interested in the regime in which the growth rates of perturbations are considerably faster. This simplifying assumption also enables the use of Floquet theory for the linear stability analysis below. Substituting $\bm{u}^{\prime}=\bm{U}^{\prime}+\bm{\widetilde{u}}^{\prime}$ and dropping quadratic terms $O(|\bm{\widetilde{u}}^{\prime}|^{2})$ gives the governing equations at first order for the perturbations,
@@
$\widetilde{f}^{\prime}=\mathrm{Re}\left[f^{\dagger}(\phi)e^{\mathrm{i}(\alpha x^{\prime}+\beta y^{\prime})+\sigma^{\prime}t^{\prime}}\right],$ (9)
@@
where $\widetilde{f}^{\prime}$ is one of $\widetilde{u}^{\prime}$, $\widetilde{v}^{\prime}$, $\widetilde{w}^{\prime}$, $\widetilde{p}\,^{\prime}$, and the form of $f^{\dagger}$ is determined by Floquet theory. The horizontal wave vector components $\alpha$ and $\beta$ are real while the frequency $\sigma^{\prime}$ is complex. Unstable modes are those with $\mathrm{Re}[\sigma^{\prime}]>0$. Substitution of Equation 9 reduces the system to a set of coupled ordinary differential equations in $\phi$:
$(S-\mathrm{Ek}L_{2})u^{\dagger}+w^{\dagger}\frac{\mathrm{d}U^{\prime}}{\mathrm{d}\phi}-\frac{\cos\theta}{|\cos\theta|}v^{\dagger}$ $=-\mathrm{i}\alpha p^{\dagger},$ (10)
$(S-\mathrm{Ek}L_{2})v^{\dagger}+w^{\dagger}\frac{\mathrm{d}V^{\prime}}{\mathrm{d}\phi}+\frac{\cos\theta}{|\cos\theta|}u^{\dagger}+|\tan\theta|w^{\dagger}$ $=-\mathrm{i}\beta p^{\dagger},$ (11)
$(S-\mathrm{Ek}L_{2})w^{\dagger}-|\tan\theta|v^{\dagger}$ $=-\frac{\mathrm{d}p^{\dagger}}{\mathrm{d}\phi},$ (12)
$\mathrm{i}\alpha u^{\dagger}+\mathrm{i}\beta v^{\dagger}+\frac{\mathrm{d}w^{\dagger}}{\mathrm{d}\phi}$ $=0,$ (13)
@@
where $S=\sigma^{\prime}+\mathrm{i}\alpha U^{\prime}+\mathrm{i}\beta V^{\prime}+W^{\prime}\frac{\mathrm{d}}{\mathrm{d}\phi}$, $L_{2}=-\alpha^{2}-\beta^{2}+\frac{\mathrm{d}^{2}}{\mathrm{d}\phi^{2}}$, $0\leq\theta\leq\pi$, and the last equation is the incompressibility constraint. The coefficients $U^{\prime}$ and $V^{\prime}$ are $2\pi$-periodic, so Floquet theory implies that $f^{\dagger}$ takes the form
$f^{\dagger}(\phi)=e^{\mathrm{i}\gamma\phi}\hat{f}(\phi)=e^{\mathrm{i}\gamma\phi}\sum_{n=-\infty}^{n=\infty}f_{n}e^{\mathrm{i}n\phi},$ (14)
@@
where $\hat{f}(\phi)$ is a $2\pi$-periodic function that is expanded in Fourier modes with complex coefficients $f_{n}$. Floquet theory allows $f^{\dagger}$ to be aperiodic since $\gamma$ need not take an integer value. However, $\gamma$ must be real for physical solutions that are bounded along $\phi$. Substitution of the Floquet ansatz (Equations 9 and 14) and the base flow (Equation 7) into the governing equations for the perturbations (Equation 8) reduces the problem to a set of coupled algebraic equations:
@@
$(\sigma^{\prime}-\mathrm{Ek}L_{2n})u_{n}+\frac{\mathrm{i}\alpha A^{\prime}}{2}(u_{n+1}+u_{n-1})-\frac{\beta sA^{\prime}}{2}(u_{n+1}-u_{n-1})$- $\frac{\mathrm{i}s\cos\theta}{|\cos\theta|}(\gamma+n)u_{n}-\frac{\mathrm{i}A^{\prime}}{2}(w_{n+1}-w_{n-1})$
$-\frac{\cos\theta}{|\cos\theta|}v_{n}=-\mathrm{i}\alpha p_{n},$ (15)
@@
$(\sigma^{\prime}-\mathrm{Ek}L_{2n})v_{n}+\frac{\mathrm{i}\alpha A^{\prime}}{2}(v_{n+1}+v_{n-1})-\frac{\beta sA^{\prime}}{2}(v_{n+1}-v_{n-1})$- $\frac{\mathrm{i}s\cos\theta}{|\cos\theta|}(\gamma+n)v_{n}+\frac{sA^{\prime}}{2}(w_{n+1}+w_{n-1})$
$+\frac{\cos\theta}{|\cos\theta|}u_{n}+|\tan\theta|w_{n}=-\mathrm{i}\beta p_{n},$ (16)
@@
$(\sigma^{\prime}-\mathrm{Ek}L_{2n})w_{n}+\frac{\mathrm{i}\alpha A^{\prime}}{2}(w_{n+1}+w_{n-1})-\frac{\beta sA^{\prime}}{2}(w_{n+1}-w_{n-1})$- $\frac{\mathrm{i}s\cos\theta}{|\cos\theta|}(\gamma+n)w_{n}-|\tan\theta|v_{n}=-\mathrm{i}(\gamma+n)p_{n},$ (17)
$\mathrm{i}\alpha u_{n}+\mathrm{i}\beta v_{n}+\mathrm{i}(\gamma+n)w_{n}$ $=0,$ (18)
@@
where $L_{2n}\equiv-\alpha^{2}-\beta^{2}-(\gamma+n)^{2}$. The linear stability of the wave can now be determined by solving this algebraic eigenvalue problem for the complex frequency $\sigma^{\prime}$ as a function of the parameters of the primary wave $\{A^{\prime},\theta,s\}$, the horizontal wave vectors of the perturbations in the comoving frame $\{\alpha,\;\beta\}$, and the Floquet parameter $\gamma$. For completeness, the energy conservation equation for this Floquet system is derived in Appendix A. Note that this equation set is similar to the analogous Floquet system for IGWs [47], but with different linear restoring terms (buoyancy versus Coriolis) and a non-zero $\hat{y}$ component of the base flow for IWs which is not present for IGWs.
@@
The eigenvalue problem given by Equations II.118 has several symmetries that reduce the parameter space over which to search for unstable modes. Several symmetries are shared with the Floquet system for IGWs [47], which has three symmetries: 1) symmetry for eigenvectors $\bm{V}_{n}\equiv[u_{n},v_{n},w_{n},p_{n}]\to\bm{V}_{n}^{*}$ (where * denotes the complex conjugate and $\bm{V}^{*}_{n}=\bm{V}_{-n}$), $(\alpha,\beta,\gamma)\to(-\alpha,-\beta,-\gamma)$, and $\sigma^{\prime}\to{\sigma^{\prime}}^{*}$, 2) invariance of $\sigma^{\prime}$ to integer shifts in the Floquet parameter $\gamma\rightarrow\gamma+1$, and 3) invariance of $\sigma^{\prime}$ in the $\beta$-plane, $\beta\rightarrow-\beta$ and $v_{n}\rightarrow-v_{n}$. The Floquet system for IWs satisfies the first two symmetries, but lacks the third because the non-zero $V^{\prime}\neq 0$ component of the base flow interacts with perturbations that have $\beta\neq 0$ (unlike for IGWs where $V^{\prime}=0$). As a result, the Floquet system for IWs must be explored in the upper (or lower) half of the $\alpha-\beta$ plane, and for all $0\leq\gamma<1$.